OpenStudy (anonymous):

1. What is the solution of n² – 49 = 0? (1 point)

OpenStudy (anonymous):

+7,-7

OpenStudy (anonymous):

A.–7 B.7 C.±7 those the choices they gave me to choose D.no solution

OpenStudy (anonymous):

Okay, you need to factor first. (n+7)(n-7) = 0 If you set n+7=0, you'll get n=-7 (-7+7)(-7-7) = 0 0(-14) = 0 Now set n-7=0 n=7 (7+7)*(7-7) = 0 14*(0) = 0 So your solutions are n=-7, 7

OpenStudy (anonymous):

so the answer will be a. -7 ?

OpenStudy (anonymous):

No, C. +/-7

OpenStudy (anonymous):

Really look over my explanation and it will make sense.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Just think about it this way... zero * (some crap) = 0, and (some crap)*zero = 0. That's why there are 2 solutions in this case.